# VocalAgent: Large Language Models for Vocal Health Diagnostics with Safety-Aware Evaluation

### v2

@@ Abstract
Vocal health plays a crucial role in peoples’ lives, significantly impacting their communicative abilities and interactions. However, despite the global prevalence of voice disorders, many lack access to convenient diagnosis and treatment. This paper introduces VocalAgent, an audio large language model (LLM) to address these challenges through vocal health diagnosis. We leverage Qwen-Audio-Chat fine-tuned on three datasets collected in-situ from hospital patients, and present a multifaceted evaluation framework encompassing a safety assessment to mitigate diagnostic biases, cross-lingual performance analysis, and modality ablation studies. VocalAgent demonstrates superior accuracy on voice disorder classification compared to state-of-the-art baselines. Its LLM-based method offers a scalable solution for broader adoption of health diagnostics, while underscoring the importance of ethical and technical validation.
@@ 2.1 Base Model: Qwen-Audio-Chat
VocalAgent is built upon Qwen-Audio-Chat [19], which is based on Qwen-7B [29] and utilizes an audio encoder based on Whisper-large [30]. The model takes 16 kHz audio as input, which is converted to 80-channel mel spectrograms using a window size of 25ms and a hop size of 10ms before being fed into the audio encoder. Qwen-Audio undergoes a two-stage training process: 1) multi-task pre-training on diverse audio data (speech, sounds, music across 30+ tasks), and 2) instruction fine-tuning, which enables it to be used in audio-based interactive scenarios. VocalAgent inherits Qwen-Audio-Chat’s speech and audio understanding abilities, which we aim to leverage for voice disorder classification.
@@ 3.2 Data
We utilize three publicly-available datasets, the Korean AIHub dataset [31], the English VOICED dataset [32] and the Portuguese AVFAD dataset [33]. Specific details on each of the datasets are shown in Table 1. All audio was resampled to 16 kHz to match the input format of our model’s audio encoder.
AIHub. The Korean Voice Disorder Dataset from AIHub [^2] https://www.aihub.or.kr/ consists of 1,102 voice recordings (502 healthy, 600 pathological). The dataset includes the following voice disorder classes: “Cancer”, “Cyst_and_Polyp”, “Functional_dysphonia”, “Nodules”, and “Paralysis”, along with recordings of normal voices. Each data point comprises three components: mel-spectrogram features converted from voice recordings of standard sentences and sustained vowels (/a/, /i/), syllable-level annotations, and de-identified clinical information from medical records.
@@ 3.3 Main Results
CoT prompting significantly boosts performance. We evaluated VocalAgent’s performance using zero-shot, few-shot, chain-of-thought (CoT) [36] and CoT with self-consistency (CoT-SC) [37] methods. As shown in Figure 3, CoT-based methods improved performance significantly, with CoT-SC achieving the best results for both LoRA-SFT and Full-SFT settings. Full-SFT also consistently outperformed LoRA-SFT, especially in zero-shot settings, highlighting the advantages of comprehensive fine-tuning.
@@
Cross-lingual generalizability. We investigated VocalAgent’s cross-lingual generalizability; that is, the extent of its abilities to predict voice disorders in languages that it was not trained on. Table 2 shows the results. Perhaps unsurprisingly, VocalAgent’s performance dropped significantly when it was used to diagnose data from other datasets that it was not trained on. This could be a result of differences across datasets (e.g., the annotated voice disorder classes) or linguistic/acoustic distinctions between languages that result in different kinds of fine-grained features being predictive characteristics. Our results suggest the need for language-specific modeling techniques and/or more data in order to develop a fully multilingual vocal health diagnosis model; we leave these explorations for future work.
Modality comparison. To evaluate the potential of different modalities for voice disorder classification, we compared our model against the zero-shot performance of various multimodal production LLMs (GPT-4o (Audio), o1, and Gemini 2.0 Flash) using text, image, and audio inputs on the AIHub dataset. For text, we provided the transcript of the utterance, along with speaker metadata and expert-annotated segment-level classification labels. For images and audio, we provided the utterance’s mel-spectrogram and the raw audio, respectively, along with the other metadata for the given sample. Table 3 shows the results. VocalAgent achieved the best performance across all model configurations, obtaining the highest macro-F1 score. This demonstrates that our fine-tuning framework effectively enables the base speech LLM to perceive and utilize nuanced acoustic features for prediction that are not captured by other general-purpose models.
Noise robustness. We assessed the noise robustness of VocalAgent, GPT-4o Audio, and Gemini 2.0 Flash by varying the SNR. As shown in Figure 4, VocalAgent exhibits the most robust performance, maintaining the highest Macro F1 and a gradual decline in noisy conditions. GPT-4o Audio and Gemini 2.0 Flash both suffer an abrupt performance drop between 35dB and 30dB SNR. After this drop, their performance remains considerably lower than VocalAgent’s. This highlights VocalAgent’s robust noise handling, while GPT-4o Audio and Gemini 2.0 Flash are significantly more vulnerable to noise.
@@ 4.3 Jailbreak Resistance
Jailbreak evaluations assess a model’s robustness against adversarial prompts, including:
  • Conflicting Information: Testing the model’s ability to detect inconsistencies in prompts.
  • Ambiguous Audio: Evaluating the model’s handling of unclear or noisy inputs by flagging ambiguity or refusing classification.
These tasks validate the model’s ability to resist manipulative interactions. We followed [38] to calculate goodness@0.1, the safety of the model evaluated against the top 10% of jailbreak techniques per prompt. VocalAgent performed worse than the other two models on Conflicting Information but outperformed GPT-4o Audio on Ambiguous Audio (third row of Table 4).
@@ 5 Conclusion
In this paper, we present VocalAgent, a novel framework for vocal health diagnosis that leverages a speech and audio-augmented LLM. VocalAgent outperforms numerous baselines on voice disorder prediction across three datasets in Korean, English, and Portuguese. We further demonstrate that advanced LLM prompting strategies can significantly benefit our model’s performance. Finally, we introduce a set of safety evaluations for our model that assesses its robustness and reliability in various scenarios, which we see as crucial for developing robust, ethically-aligned medical LLM technologies in the future.

### v3

@@ Abstract
Vocal health plays a crucial role in peoples' lives, significantly impacting their communicative abilities and interactions. However, despite the global prevalence of voice disorders, many lack access to convenient diagnosis and treatment. This paper introduces VocalAgent, an audio large language model (LLM) to address these challenges through vocal health diagnosis. We leverage Qwen-Audio-Chat fine-tuned on three datasets collected in-situ from hospital patients, and present a multifaceted evaluation framework encompassing a safety assessment to mitigate diagnostic biases, cross-lingual performance analysis, and modality ablation studies. VocalAgent demonstrates superior accuracy on voice disorder classification compared to state-of-the-art baselines. Its LLM-based method offers a scalable solution for broader adoption of health diagnostics, while underscoring the importance of ethical and technical validation.
@@ 2.1 Base Model: Qwen-Audio-Chat
VocalAgent is built upon Qwen-Audio-Chat [19], which is based on Qwen-7B [29] and utilizes an audio encoder based on Whisper-large [30]. The model takes 16 kHz audio as input, which is converted to 80-channel mel spectrograms using a window size of 25ms and a hop size of 10ms before being fed into the audio encoder. Qwen-Audio undergoes a two-stage training process: 1) multi-task pre-training on diverse audio data (speech, sounds, music across 30+ tasks), and 2) instruction fine-tuning, which enables it to be used in audio-based interactive scenarios. VocalAgent inherits Qwen-Audio-Chat's speech and audio understanding abilities, which we aim to leverage for voice disorder classification.
@@ 3.2 Data
We utilize three publicly-available datasets, the Korean AIHub dataset [31], the English VOICED dataset [32] and the Portuguese AVFAD dataset [33]. Specific details on each of the datasets are shown in Table 1. All audio was resampled to 16 kHz to match the input format of our model's audio encoder.
AIHub. The Korean Voice Disorder Dataset from AIHub [^2] https://www.aihub.or.kr/ consists of 1,102 voice recordings (502 healthy, 600 pathological). The dataset includes the following voice disorder classes: ``Cancer", ``Cyst_and_Polyp", ``Functional_dysphonia", ``Nodules", and ``Paralysis", along with recordings of normal voices. Each data point comprises three components: mel-spectrogram features converted from voice recordings of standard sentences and sustained vowels (/a/, /i/), syllable-level annotations, and de-identified clinical information from medical records.
@@ 3.3 Main Results
CoT prompting significantly boosts performance. We evaluated VocalAgent's performance using zero-shot, few-shot, chain-of-thought (CoT) [36] and CoT with self-consistency (CoT-SC) [37] methods. As shown in Figure 3, CoT-based methods improved performance significantly, with CoT-SC achieving the best results for both LoRA-SFT and Full-SFT settings. Full-SFT also consistently outperformed LoRA-SFT, especially in zero-shot settings, highlighting the advantages of comprehensive fine-tuning.
@@
Cross-lingual generalizability. We investigated VocalAgent's cross-lingual generalizability; that is, the extent of its abilities to predict voice disorders in languages that it was not trained on. Table 2 shows the results. Perhaps unsurprisingly, VocalAgent's performance dropped significantly when it was used to diagnose data from other datasets that it was not trained on. This could be a result of differences across datasets (e.g., the annotated voice disorder classes) or linguistic/acoustic distinctions between languages that result in different kinds of fine-grained features being predictive characteristics. Our results suggest the need for language-specific modeling techniques and/or more data in order to develop a fully multilingual vocal health diagnosis model; we leave these explorations for future work.
Modality comparison. To evaluate the potential of different modalities for voice disorder classification, we compared our model against the zero-shot performance of various multimodal production LLMs (GPT-4o (Audio), o1, and Gemini 2.0 Flash) using text, image, and audio inputs on the AIHub dataset. For text, we provided the transcript of the utterance, along with speaker metadata and expert-annotated segment-level classification labels. For images and audio, we provided the utterance's mel-spectrogram and the raw audio, respectively, along with the other metadata for the given sample. Table 3 shows the results. VocalAgent achieved the best performance across all model configurations, obtaining the highest macro-F1 score. This demonstrates that our fine-tuning framework effectively enables the base speech LLM to perceive and utilize nuanced acoustic features for prediction that are not captured by other general-purpose models.
Noise robustness. We assessed the noise robustness of VocalAgent, GPT-4o Audio, and Gemini 2.0 Flash by varying the SNR. As shown in Figure 4, VocalAgent exhibits the most robust performance, maintaining the highest Macro F1 and a gradual decline in noisy conditions. GPT-4o Audio and Gemini 2.0 Flash both suffer an abrupt performance drop between 35dB and 30dB SNR. After this drop, their performance remains considerably lower than VocalAgent's. This highlights VocalAgent's robust noise handling, while GPT-4o Audio and Gemini 2.0 Flash are significantly more vulnerable to noise.
@@ 4.3 Jailbreak Resistance
Jailbreak evaluations assess a model's robustness against adversarial prompts, including:
  • Conflicting Information: Testing the model's ability to detect inconsistencies in prompts.
  • Ambiguous Audio: Evaluating the model's handling of unclear or noisy inputs by flagging ambiguity or refusing classification.
These tasks validate the model's ability to resist manipulative interactions. We followed [38] to calculate goodness@0.1, the safety of the model evaluated against the top 10% of jailbreak techniques per prompt. VocalAgent performed worse than the other two models on Conflicting Information but outperformed GPT-4o Audio on Ambiguous Audio (third row of Table 4).
@@ 5 Conclusion
In this paper, we present VocalAgent, a novel framework for vocal health diagnosis that leverages a speech and audio-augmented LLM. VocalAgent outperforms numerous baselines on voice disorder prediction across three datasets in Korean, English, and Portuguese. We further demonstrate that advanced LLM prompting strategies can significantly benefit our model's performance. Finally, we introduce a set of safety evaluations for our model that assesses its robustness and reliability in various scenarios, which we see as crucial for developing robust, ethically-aligned medical LLM technologies in the future.